### v2

@@ Abstract
In conclusion, we introduced ProteinGuide, a broadly applicable framework that enables conditioning of pre-trained protein generative models on any property of interest that can be encoded in a predictive model, without having to re-train the pre-trained model. An important application enabled by our method is to incorporate feedback from experimental measurements into the generative process.
@@ Acknowledgments
Methods
Background on discrete state-space diffusion and flow models
While diffusion models were initially developed for continuous state-spaces [13, 15, 29], recent work has extended these frameworks to discrete domains such as text, molecular graphs, and protein sequences. These extensions enable generative modeling of inherently discrete data without requiring continuous relaxations. There are a number of works defining discrete-time diffusion processes in discrete spaces [19, 30]. They were later extended to continuous-time [31, 32, 33] leveraging the formulation of continuous-time Markov chains (CTMCs), which describe stochastic processes where a system transitions between discrete states over continuous time. The dynamics of a CTMC are fully determined by an initial distribution $p(x_{0})$ and transition rates $R_{t}(x,\tilde{x})$ that specify the instantaneous probability of transitioning from state $x$ to state $\tilde{x}$ at time $t$:
$p(x_{t+dt}=\tilde{x}|x_{t}=x)=\delta_{x,\tilde{x}}+R_{t}(x,\tilde{x})dt+O(dt^{1+\epsilon}).$ (1)
@@ Methods
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p_{\theta}(x_{1}|x_{t},t)\right],$ (2)
@@
$R^{\theta}_{t}(x_{t},x_{t+dt})=\mathbb{E}_{x_{1}\sim p_{\theta}(x_{1}|x_{t})}\Big{[}\overrightarrow{R_{t}}(x_{t+dt},x_{t}|x_{1})\Big{]},$ (3)
@@
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right].$ (4)
@@
$R^{\theta,i}_{t}(x_{t},x_{t+dt}^{i})=\mathbb{E}_{x_{1}^{i}\sim p_{\theta}(x_{1}^{i}|x_{t},t)}\Big{[}\overrightarrow{R_{t}^{i}}(x_{t+dt}^{i},x_{t}^{i}|x_{1}^{i})\Big{]},$ (5)
@@
A popular instantiation of both discrete state-spaces diffusion and flow models uses masking as the per-dimension noising process, where tokens in a sequence are progressively replaced with mask tokens during the forward process and the model learns to recover the original tokens during the reverse process. For example, the masking conditional flow can be defined as $p^{\text{mask}}(x_{t}^{i}|x_{1}^{i})=\text{Cat}(t\delta\{x_{1}^{i},x_{t}^{i}\}+(1-t)\delta\{?,x_{t}^{i}\})$, where $\delta\{i,j\}$ is the Kronecker delta which is $1$ when $i=j$, 00 otherwise, and “?” denotes a mask token. Campbell et al. [37] showed that under the masking noise process, the unconditional rate matrix used in sampling defined in Equation 5 simplifies to
$R^{\theta,i}_{t}(x_{t},x_{t+dt}^{i})=\frac{p_{\theta}(x_{1}^{i}\,|\,x_{t},t)}{1-t}\delta\{x_{t}^{i},?\}.$ (6)
@@
This formulation has connections to Masked Language Models and order-agnostic autoregressive models, which we will describe in detail next.
Unifying connections between generative models on discrete state-spaces
Connections among masked discrete diffusion models, Masked Language Models (MLM) and order-agnostic autoregressive models (OA-AR) were initially pointed out by Austin et al. [19] and Hoogeboom et al. [20]. Recent works [21, 22, 23, 37, 38] further clarified these connections, including the connections to CTMC-based formulation of diffusion and flow models [31, 33, 37], and showed that the continuous-time ELBO used to train masked diffusion models is essentially a weighted sum of cross entropy losses predicting the masked states from unmasked positions, where the proportion of masked positions ranges from 0 (complete masking) to 1 (no masking). For particular choices of the noise schedules, these training objectives can be shown to be equivalent to the training objective of order-agnostic autoregressive models [42], (generative) Masked Language Models [19], or discrete flow-matching models [37, 38] under a masking process. Since ESM3 [4] and ProteinMPNN [1] were trained under the MLM and OA-AR objectives respectively, we can view them as discrete state-space diffusion and flow models, and sample from them as such, leveraging the generative rate matrices from Equation 3 induced by the trained denoising model, $p_{\theta}(x_{1}|x_{t})$. This formulation not only provides a principled way to sample from these models, but also offers additional sampling flexibility, including, but not limited to, the usage of different noise schedules during sampling from those used during training, and corrector-sampling [37, 38], which have been shown to sometimes improve sampling results compared to prior heuristics [43] for sampling from generative Masked Language Models [38]. In addition, this formulation unlocks the capacity to condition the generative process via guidance, which we introduce in a later section.
Here, we provide a self-contained derivation of the equivalence between the loss functions of discrete flow matching, Masked Language Models, and order-agnostic autoregressive models.
Equivalence between flow matching on discrete state-space and masked language modeling
Recall from Equation 4 that the flow matching loss is
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right].$ (7)
@@
Our goal is to demonstrate that this loss function is equivalent to the loss function used by a class of Masked Language Models (MLMs), that we will refer to as “generative Masked Language Models”. The key distinction between generative Masked Language Models and their non-generative counterparts, is that during training, generative MLMs mask tokens with a masking probability, $t\in[0,1]$, sampled from a distribution $p(t)$ that has full support on the interval $[0,1]$. This requirement is met by MLMs such as ESM3 [4], but, importantly, is not met by BERT-style MLMs such as ESM2 [44], since these models use a fixed percentage of tokens that are masked at each step. Letting $x_{t}$ refer to the masked sequence, we have that the generative masked language modeling objective function is
@@
$\mathcal{L}_{\text{MLM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right],$ (8)

### v4

@@ Discussion
We introduced ProteinGuide, a broadly applicable framework that enables conditioning of pre-trained protein generative models on any property of interest that can be encoded in a predictive model, without having to re-train the pre-trained model. An important application enabled by ProteinGuide is to nimbly incorporate feedback from experimental measurements into the generative process, on-the fly. Through in silico experiments in three distinct design settings—interpolative, extrapolative, and multi-property Pareto-extrapolative—we demonstrated the utility of ProteinGuide in a variety of application settings of interest to protein engineers, across a variety of pre-trained generative models, and spanning numerous protein properties. In our in vivo experiment to engineer an adenine base editor with improved activity, we demonstrated that a single round of ProteinGuide yielded variants with activity higher than that produced by seven rounds of directed evolution.
Guidance, and hence ProteinGuide, differs from the common practices of post hoc filtering and fine-tuning in a few important ways. First, it uses a separate predictive model to modulate the sampling process from a pre-trained model, without having to train/re-train any generative model. Second, it uses the principle of Bayes’ rule to statistically, coherently blend information from the pre-trained model with the predictive model, thereby enabling generation of variants that the pre-trained model would not otherwise produce. Moreover, the coherent blending mitigates worries about “forgetting” that often emerge from fine-tuning.
Beyond straightforward application of our on-the-fly guidance approach, one might consider combining ProteinGuide with other Bayesian-inspired frameworks that involve re-training a pre-trained model towards a reward-weighted posterior distribution [117, 89, 120, 18], particularly in cases when the targeted region is significantly out of distribution for the pre-trained generative model. Pure guidance might be challenging in such a situation. Separately, it might prove promising to combine ProteinGuide with MCMC-based methods, such as recent works for guiding generative models with Sequential Monte Carlo (SMC) [122, 67, 99], to improve the generation quality, albeit at the expense of added sampling computation time.
We anticipate future work will further extend and build upon our work. One interesting line of exploration might involve applying ProteinGuide in conjunction with guidance methods for continuous spaces to guide multimodal generative models of both structure and sequence [13, 60].
@@ Acknowledgments
Supplementary Materials for
ProteinGuide: On-the-fly property guidance for protein sequence generative models
Junhao Xiong, Ishan Gaur, Maria Lukarska, Hunter Nisonoff,
Luke M. Oltrogge, David F. Savage, Jennifer Listgarten
Corresponding authors. Email: savage@berkeley.edu, jennl@berkeley.edu
These authors contributed equally to this work.
This PDF file includes:
Materials and Methods
Supplementary Text
Supplementary Tables
Supplementary Figures
1 Materials and Methods
The Materials and Methods are organized as follows. We first provide an overview of the technical underpinning of ProteinGuide (Section 1.1), followed by more detailed technical background (Section 1.2). We proceed to show the equivalence of training objectives (Section 1.3) and sampling algorithms (Section 1.4) for a broad class of models, including masking-based diffusion/flow-matching models (MDMs/MFMs), generative Masked Language Models (MLMs), and any-order autoregressive models (AO-ARMs). The sections afterwards provide full experimental details for all experiments, including an overview for the in silico experiments (Section 1.5), followed by dedicated sections for each experiment (Section 1.6-1.11).
1.1 Overview of ProteinGuide
Conceptually, the problem of guidance involves taking a pre-trained generative model, $p(x)$, and a classifier- or regression-derived likelihood, $p(y|x)$, to obtain $p(x|y)=p(y|x)p(x)/p(y)$, without re-training the underlying generative model. At bottom, designing guidance strategies amounts to exploiting local structure in the generative process so that this reweighting can be implemented tractably and efficiently during generation, while producing samples from the correct conditional distribution at the end of the process. As an example, diffusion models on continuous state-spaces generate data by learning to reverse a “noising” process, most commonly instantiated as gradually corrupting the data towards pure Gaussian noise [100, 101, 46]. Intuitively, guidance in diffusion models is achieved by leveraging a key property of the underlying generative process—namely, that the model learns to map from a simple distribution (e.g., a standard normal distribution) to a complex data distribution through a series of infinitesimal steps. Each step “points” towards directions with higher likelihood under the data distribution, and can be “nudged”, at generation time, towards directions that better satisfy the user-specified likelihood, so that the process ends with samples from the desired conditionals. These infinitesimal steps are formalized as gradients of noise-dependent distributions, known as score functions, and Bayes’ Rule can be simplified considerably when considering only these gradients, enabling guidance for these models [100, 101, 26].
For discrete-space generative models, however, these gradients are not defined, and such an approach cannot be readily adopted. Nevertheless, discrete-space diffusion and flow-based models also admit a natural local structure, but in a different form. These models learn to reverse a user-defined discrete noising process, such as introducing random mutations, or accumulating mask tokens [4, 49]. Such process can be formalized with continuous-time Markov chains (CTMCs), which can be described as a generative process wherein a state remains unchanged over some amount of holding time, then changes (jumps) to a new state. In practice, these two processes of holding, and jumping, are all jointly encoded in an object known as the (time-dependent) rate matrix, which describes the time derivative of the probability of transitioning from any state to any other state at a given time. These rate matrices play a similar role to the score functions in continuous-space diffusion models, and sampling corresponds to integrating the CTMCs, for example, by straightforward Euler integration [13] or with more involved methods [37, 38, 12]. To make learning tractable, these models typically assume that the noising processes in each dimension are independent of one another (but not necessarily identical), with the important consequence that at any given time in the continuous-time noising process, the probability that two or more dimensions of the current state jump (transition) at exactly the same time is zero [12, 13]. Consequently, at any given time, the number of entries in the rate matrix that are non-zero is linear, rather than exponential, in the input space, making such models tractable.
It turns out that the same assumption which makes training these models tractable also makes guidance tractable for them. By reformulating Bayes’ Rules in terms of the rates, we find that the conditional rates at each sampling step can be obtained by modulating the unguided rates with a likelihood ratio describing how much the guidance signal increases (or decreases) in an infinitesimal time step. Since at any given time step, only transitions with a single mutation have non-zero rates, one can tractably compute the normalizing constant for any given time step [82]. Since this procedure affects only sampling and not training, one can sample from any desired conditional distribution for which a user-specified likelihood is available, without having to re-train the underlying pre-trained generative model. We refer to this work as Discrete Guidance and the reader to Nisonoff et al. [82] for more details.
While the framework just presented applies generally to discrete-space diffusion and flow-based models under a broad set of noising processes, one popular instantiation worth considering is a noising process based on masking, where tokens in a sequence are progressively replaced with mask tokens during the noising process and the model learns to recover the original tokens during the denoising process. These models share deep connections to two other popular classes of discrete-space generative models, Masked Language Models (MLMs) with masking rates covering the range $0\%-100\%$,, sometimes known as generative MLMs, and any-order autoregressive models (AO-ARMs). Although conceptually masked diffusion/flow-matching models (MDMs/MFMs), MLMs, and AO-ARMs seem to be quite different, it turns out that the loss functions used for training them are equivalent. While connections between these training objectives has been previously noted [4, 49, 98, 94, 83], the practical utility of this observation has been unclear. Herein, we show that this connection enables a general framework for guiding all of these model classes. Moreover, we additionally show for the first time that sampling from a masked diffusion/flow-matching model can be accomplished more simply, yet exactly, by sampling from a corresponding, trivial-to-construct AO-ARM, which leads to a more efficient, exact guidance algorithm for MDMs, MFMs, MLMs, and AO-ARMs. Altogether, our unifying view of training and sampling of all four of these model classes under a single framework, enables ProteinGuide to guide a broad class of discrete-space generative models, making it readily applicable to many popular protein sequence generative models commonly used by practitioners.
1.2 Technical background
Here we give a more mathematical treatment of the overview above, providing the technical background important for understanding the results shown in later sections.
1.2.1 Discrete-space diffusion and flow models (DDMs and DFMs)
[…]
$p(X_{t+dt}=\tilde{x}|X_{t}=x)=\delta_{x,\tilde{x}}+R_{t}(x,\tilde{x})dt+O(dt^{1+\epsilon}).$ (S1)
@@ 1.2.1 Discrete-space diffusion and flow models (DDMs and DFMs)
$\mathcal{L}_{\text{FM}}=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),\,t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log p_{\theta}(x_{1}|x_{t},t)\right],$ (S2)
@@
$R^{\theta}_{t}(x_{t},x_{t+dt})=\mathbb{E}_{x_{1}\sim p_{\theta}(x_{1}|x_{t})}\Big[\overrightarrow{R_{t}}(x_{t+dt},x_{t}|x_{1})\Big],$ (S3)
@@
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right].$ (S4)
@@
$R^{\theta,i}_{t}(x_{t},x_{t+dt}^{i})=\mathbb{E}_{x_{1}^{i}\sim p_{\theta}(x_{1}^{i}|x_{t},t)}\Big[\overrightarrow{R_{t}^{i}}(x_{t+dt}^{i},x_{t}^{i}|x_{1}^{i})\Big],$ (S5)
@@
A popular instantiation of both discrete state-spaces diffusion and flow models uses masking as the per-dimension noising process, where tokens in a sequence are progressively replaced with mask tokens during the forward process and the model learns to recover the original tokens during the reverse process. For example, the masking conditional flow can be defined as $p^{\text{mask}}(x_{t}^{i}|x_{1}^{i})=\text{Cat}(t\delta\{x_{1}^{i},x_{t}^{i}\}+(1-t)\delta\{?,x_{t}^{i}\})$, where $\delta\{i,j\}$ is the Kronecker delta which is $1$ when $i=j$, 0 otherwise, and “?” denotes a mask token. Campbell et al. [13] showed that under the masking noise process, the unconditional rate matrix used in sampling defined in Equation (S5) simplifies to
$R^{\theta,i}_{t}(x_{t},x_{t+dt}^{i})=\frac{p_{\theta}(x_{1}^{i}\,|\,x_{t},t)}{1-t}\delta\{x_{t}^{i},?\}.$ (S6)
@@
This formulation has connections to Masked Language Models and any-order autoregressive model, which we will describe in detail next.
1.2.2 Masked language models (MLMs)
@@
Masked Language Models (MLMs) loosely refer to a broad class of models that involves training a model to recover sequences of discrete tokens given the input where some of the tokens have been replaced by a special token known as the “mask” token. Herein we focus on a class of Masked Language Models (MLMs), that we will refer to as “generative Masked Language Models” [36, 14]. The key distinction between generative Masked Language Models and their non-generative counterparts [25], is that during training, generative MLMs mask tokens with a masking probability, $t\in[0,1]$, sampled from a distribution $p(t)$ that has full support on the interval $[0,1]$. This requirement is met by MLMs such as ESM3 [44], but, importantly, is not met by BERT-style MLMs such as ESM2 [69], since these models use a fixed percentage of tokens that are masked at each step. Letting $x_{t}$ refer to the masked sequence with masking proportion denoted by $t$, we have that the generative masked language modeling objective function is
@@
$\mathcal{L}_{\text{MLM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right],$ (S7)
There are various approaches that are used to sample from a generative MLM. Some common approaches include directly decoding all the tokens from the fully masked sequence, or decoding the sequence iteratively [44]. For iteratively decoding, a commonly used heuristic to choose the decoding order which prioritizes the token to be unmasked according to some confidence [14].
1.2.3 Any-order autoregressive models (AO-ARMs)
Any-order autoregressive models (AO-ARMs) differ from standard autoregressive models by training on arbitrary permutations of the sequence ordering [112, 49]. Consider a sequence $x=(x^{1},x^{2},...,x^{D})$ from our data distribution. In an any-order autoregressive model, we sample a permutation $\sigma$ of indices $\{1,2,...,D\}$ and factorize the joint distribution as:
$p_{\theta}(x)=\prod_{i=1}^{D}p_{\theta}(x^{\sigma(i)}|x_{1}^{\sigma(<i)})$ (S8)
where $x^{\sigma(<i)}$ denotes $(x^{\sigma(1)},x^{\sigma(2)},...,x^{\sigma(i-1)})$. The AO-ARM objective is to maximize the expected log-likelihood over all possible permutations:
$\mathcal{L}_{\text{AO-ARM}}$ $=\mathbb{E}_{x\sim p_{\text{data}}(x),\sigma\sim\text{Unif}(\text{Perm}(D))}\left[\sum_{i=1}^{D}\log p_{\theta}(x^{\sigma(i)}|x^{\sigma(<i)})\right]$ (S9)
Sampling from an AO-ARM [112, 127, 106, 49] involves iteratively constructing a sample, $x=(x^{1},\ldots,x^{D})$ of length $D$, starting from a fully masked sequence, by first sampling an ordering of positions uniformly at random, then sample iteratively from the conditional distribution of the model given this ordering.
1.2.4 Conditional generation with guidance
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (S10)
$\nabla_{x_{t}}\log p^{(\gamma)}_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(x_{t})+\gamma\nabla_{x_{t}}\log p_{t}(y|x_{t})$ (S11)
$p(x_{t+dt}|x_{t},y)=\frac{p(y|x_{t+dt},x_{t})p(x_{t+dt}|x_{t})}{\sum\limits_{x^{\prime}_{t+dt}}p(y|x^{\prime}_{t+dt},x_{t})p(x^{\prime}_{t+dt}|x_{t})},$ (S12)
where $x_{t+dt}$ and $x_{t}$ are separated by a infinitesimal time step $dt$, and $y$ is the desired property we wish to condition on. In a $D$-dimensional discrete state-space where each dimension has cardinality $S$, there are $S^{D}$ terms in the normalizing constant of Equation S12, rendering it generally intractable. Each term arises from the fact that, without constraints, any state can be reached from any other state. Intuitively, tractability can only be achieved if the number of terms is not exponential, thereby implying some constraints on the allowed state transitions. Next, we shall see how the continuous-time formulation of discrete state-space diffusion and flow models enables us to tractably compute the normalizing constant by imposing such constraints, without losing any model expressibility.
1.2.5 Guidance in discrete state-spaces
[…]
or derive predictor-free guidance discrete state-spaces:
$R_{t}^{(\gamma)}(x,\tilde{x}|y)=R_{t}(x,\tilde{x}|y)^{\gamma}R_{t}(x,\tilde{x})^{1-\gamma}$ (S15)
$\log\frac{p_{\phi}(y|\tilde{x},t)}{p_{\phi}(y|x,t)}\approx(\tilde{x}-x)^{T}\nabla_{x}\log p_{\phi}(y|x,t)$ (S16)
requiring only one forward and one backward pass of the predictor model instead of $D\times(S-1)+1$ forward passes, making estimation of the guide-adjusted rates $O(1)$ rather than $O(D\times S)$. With multiple particles, TAG can be further improved by acting as a proposal distribution in a Sequential Monte Carlo (SMC) algorithm similar to Wu et al. [122]. While there are other recently proposed approaches for guiding discrete diffusion models [67, 96, 99], ProteinGuide has the benefit of being both exact and general, applicable to both diffusion and flow-based models, making it a general and flexible framework for guiding a broad class of discrete state-space generative models (for discussion on related work, see Section 2.1).
1.3 Unified connections of training objectives for masked-based generative models
Connections among masked discrete diffusion models, Masked Language Models (MLM) and any-order autoregressive models (AO-ARMs) were initially pointed out by Austin et al. [4] and Hoogeboom et al. [49]. Recent works [98, 94, 83, 13, 32] further clarified these connections, including the connections to CTMC-based formulation of diffusion and flow models [12, 74, 13], and showed that the continuous-time ELBO used to train masked diffusion models is essentially a weighted sum of cross entropy losses predicting the masked states from unmasked positions, where the proportion of masked positions ranges from 0 (complete masking) to 1 (no masking). For particular choices of the noise schedules, these training objectives can be shown to be equivalent to the training objective of any-order autoregressive model s [112], (generative) Masked Language Models [4], and discrete flow-matching models [13, 32] under a masking process. Since ESM3 [44] and ProteinMPNN [20] were trained under the MLM and AO-ARM objectives respectively, and owing to the sampling equivalency between AO-ARMs and diffusion models, we can view ESM3 and ProteinMPNN as discrete state-space diffusion and flow models, and sample from them as such, leveraging the generative rate matrices from Equation (S3) induced by the trained denoising model, $p_{\theta}(x_{1}|x_{t})$. This formulation not only provides a principled way to sample from these models, but also offers additional sampling flexibility, including, but not limited to, the usage of different noise schedules during sampling from those used during training, and corrector-sampling [13, 32], which have been shown to sometimes improve sampling results compared to prior heuristics [14] for sampling from generative Masked Language Models [32]. In addition, this formulation unlocks the capacity to condition the generative process via guidance, which we introduce in a later section.
Here, we provide a self-contained derivation of the equivalence between the loss functions of masked flow matching (MFMs), Masked Language Models (MLMs), and any-order autoregressive models (AO-ARMs).
1.3.1 Equivalence between MFMs and MLMs
Recall from Equation (S2) that the flow matching loss is
$\mathcal{L}_{\text{FM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t},t)\right].$
Our goal is to demonstrate that this loss function is equivalent to the loss function of generative Masked Language Models (MLMs), which, recalling from Equation (S7), can be written as:
$\mathcal{L}_{\text{MLM}}$ $=\mathbb{E}_{x_{1}\sim p_{\text{data}}(x_{1}),t\sim p(t),\,x_{t}\sim p_{t|1}(x_{t}|x_{1})}\left[\log\prod_{i=1}^{D}p_{\theta}(x_{1}^{i}|x_{t})\right],$